OpenStudy (anonymous):

Just need a quick check up on my work, The center is the origin, a=8, b=6, and the major axis is parallel to the y-axis. Write the equation of the ellipse.

OpenStudy (anonymous):

Center is (0,0) so (x-0)^2/6+9y-0)^2/8=1

OpenStudy (loser66):

the denominators: a^2, b^2. simplify and you are ok

OpenStudy (anonymous):

\[\frac{ (x-0)^{2} }{ 6 }+\frac{ (y-0)^{2} }{ 8 }=1\]

OpenStudy (anonymous):

oh so it would be (x-0)^2/3 + (y-0)^2/4=1

OpenStudy (loser66):

hey, ^2, not /2 \[\dfrac{x^2}{6^2}+\dfrac{y^2}{8^2}=1\\\dfrac{x^2}{36}+\dfrac{y^2}{64}=1\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (loser66):

np